My problem is; Kristi jogged for 3/5 of an hour, swam for 1/2 of an hour, and rode her bicycle for 3/4 of an hour. How long did she exercise?
Found 2 solutions by adamchapman, mommasita:
Answer by adamchapman(301)

3/5+1/2+3/4=t
where "t" is the time she spent exercising
multiply both sides by all the denominators:
3(2*4)+1(5*4)+3(5*3)=(5*2*4)t
24+20+75=40t
t=(24+20+75)/40
t=2.975 hours
I hope this helps
P.S. I am trying to start up my own homework help website. I would be extremely grateful if you would e-mail me some feedback on the help you received to adam.chapman@student.manchester.ac.uk

Answer by mommasita(1)

3/5 + 1/2 + 3/4
3/5 = 12/20
1/2 = 10/20
3/4 = 15/20

You have to find the common denomonator which is where 20 comes in and becomes the denomonator. Then 20 divided by 5 is 4 and 4x3=12.
20 divided by 2 is 10 and 10x1=10.
20 divided by 4 is 5 and 5x3=15.
So now you can add yoru problem.
12/20 + 10/20 + 15/20 = 37/20.
Now you have to reduce this down as it is an improper fraction.
Divide 37 by 20 and you get 1 with remainder of 17.
So yoru final answer would be 1 17/20 is how long Kristi exercised.